# Scalable simulation of non-equilibrium quantum dynamics via classically optimised unitary circuits

### v1

@@ I Introduction
In this paper, we propose a tensor network (TN) approach to classically optimise brickwall (BW) unitary circuits. Our method approximates the time propagator for Hamiltonian dynamics at small times as a matrix product operator (MPO) [53, 54, 55]. This MPO is then used along with the polar decomposition [56, 46] to variationally optimise a unitary BW circuit on a classical computer by maximising their “overlap”. The advantage of using an MPO to approximate the time propagator is that it can reach a desired accuracy with resources which scale only linearly in system size [57, 58, 59]. Furthermore, we find that the optimised circuits outperform Trotterization in both accuracy and the number of two-qubit entangling layers needed to implement the dynamics. The classically optimised circuit can then be used many times on a digital quantum computer to reach the total target time.
Our paper is structured as followed. In Sec. II, we give a brief introduction to simulating quantum dynamics with Trotterization. The optimisation procedure is then explained in Sec. III. We demonstrate our approach in Sec. IV for various three-body Hamiltonians. Using numerous metrics, we compare our results to Trotterization to demonstrate the practical advantage of classically optimised BW circuits. We also benchmark convergence for our method. In Sec. V, we explain how to estimate the optimal time step which minimises the compounded error of decoherence from the quantum computer, and the variational error of the classically optimised circuit. We give our conclusions in Sec. VI.
@@ III Classical optimisation of unitary circuits
In this section, we explain how one can optimise a BW unitary circuit with two-site gates and depth $M$ to approximate $U(\Delta t)$ (for small $\Delta t$) in a way that is scalable with system size $N$. They are arranged compactly such that each layer has a unitary gate acting on each lattice site (except for lattice sites at the edges), see Figs. 2(a, b) for visualisations of circuit depths $M=3,4$. Notice that, unlike the Trotter circuits, there is no direct way to determine an optimal BW circuit for general $\hat{H}$. However, they have the distinct advantage that they can be easily implemented on current quantum computers, see App. B.
@@ III.1 MPO approximation of $U(\Delta t)$
The application of each gate will give a new MPO with a bond dimension $\chi_{\rm exact}=r\chi$, where $r>1$ is some integer which depends on the circuit. Thus, the bond dimension of the resulting MPO will grow exponentially in the number of applications of the Trotterized circuit. To prevent this, at each time step, we approximate the MPO with another MPO with a smaller bond dimension $\chi^{\prime}$ such that $\chi\leq\chi^{\prime}\leq\chi_{\rm exact}$. In practice, this is achieved by compressing the MPO using SVDs after each application of unitary gate [64]. We allow the bond dimension to grow dynamically, with its value determined through the SVD. We use an SVD truncation error of $\epsilon=10^{-16}$ to ensure that the truncation error is minimal  [^2] See Ref. [64] for a definition of the truncation error.. Notice that this set-up allows for each individual bond dimension between lattice sites to take its own value.
@@ III.4 Contracting the environment
As is often done in TN methods, it is convenient and efficient to recycle partially contracted networks [64]. For this method, it will be useful to save each of the blocks $L_{j}$ and $R_{j}$ to memory for two reasons. The first is the blocks $L_{k-1}$ and $R_{k+1}$ will be the same for each of the $M$ gates which act on site $k$. They can be calculated and then recycled to update each of the $M$ gates. The second reason is that the blocks $L_{j}$ and $R_{j}$ can be used iteratively to calculate the blocks at all sites $j$. That is, $L_{j}$ can be calculated using $L_{j-1}$ and $R_{j}$ can be calculated using $R_{j+1}$. This procedure is illustrated in Fig. 6 for a circuit depth $M=3$, although the exact calculations will differ for other circuit depths. The left column shows how to construct the blocks for even $j$, and the right column for odd $j$. The block $L_{j-1}$ is contracted with all the gates which act on both sites $j-1$ and $j$, along with the adjoint of the MPO tensor at site $j$ to form the block $L_{j}$. Similarly, the block $R_{j+1}$ is contracted with the gates which act on both sites $j$ and $j+1$, along with the adjoint of the MPO tensor at site $j$ to give $R_{j}$. The cost of expanding such blocks is at most $\mathcal{O}\left(\chi^{2}d^{2+2\lceil M/2\rceil}+\chi Md^{4+2\lceil M/2\rceil}\right)$, and storing to memory is at most $\mathcal{O}\left(\chi d^{2\lceil M/2\rceil}\right)$ per block.
@@ IV.1.3 The next-nearest-neighbour Ising model
Unlike the two previous models, we find that the optimisation for the NNNI model is highly susceptible to becoming stuck in local minima. Because of this, optimising the circuits to achieve a scaling that beats $\mathcal{O}(\Delta t)$ is extremely difficult, and only happens on rare occasions for some choice of initial guess. This effect becomes more apparent for larger system sizes. As such, this model serves as an interesting example where the solution is known explicitly, but the variational optimisation method is unable to reliably obtain a good result. Instead, we propose an initial guess which uses the structure of the circuit in Fig. 9(a), but only for the SWAP gates, see Fig. 9(b). All other gates are set to identity. With this choice of initial guess, we find the error metrics converge to those of the second order Trotterization solution, see Figs. 9(c, d) for the error density, and Figs. 9(e, f) for the conservation of energy. These results could motivate the future investigation of strategies to systemically suggest more optimal choices of initial guesses.
@@ IV.2 Number of CNOT layers
Table 1: The number of CNOT layers required to implement a single time step of the Trotterized circuits and BW circuits, and their associated error density $\delta(\Delta t)$ (rounded to three significant figures), for time step $\Delta t=0.1$ and system size $N=8$. Results are shown for Cluster Ising (CI) model, the PXP model and a next-nearest-neighbour Ising (NNNI) model. The order of the approximation, $k$, is given for each circuit. Note that the error density scales as $\mathcal{O}(\Delta t^{k+1})$ for small $\Delta t$. The CI model is for $g=-0.75$ and the NNNI model is for $g_{x}=g_{zz}=g_{z1z}=1$.
@@ V Optimal time step
We have explained how to optimise a BW circuit to efficiently simulate quantum dynamics. However, it is not immediately clear what time step $\Delta t$ should be chosen to simulate the dynamics. On one hand, we have shown that choosing a smaller time step can reduce the algorithmic error for evolving to time $t$. Indeed, if a BW circuit of depth $M$ gives a $k$th order approximation (e.g. $M=3$ for the CI model has order $k=2$), then the cost of evolving to time $t$ is $\mathcal{O}(t\Delta t^{k})$. This error is illustrated by the dashed red line in Fig. 11. This entices us to pick a small time step, resulting in a deeper quantum circuit. On the other hand, quantum computers are never completely isolated from their environment. This introduces errors which are native to the device. We assume that this error is exponential in the circuit depth, as shown by the blue dashed line in Fig. 11. This indicates that our simulations would benefit from a shallower quantum circuit. The total simulation error is a combination of the two, shown by the black curve, and the optimal time step is the one which minimises this error. In what follows, we explain how to estimate each of these errors to guide us in determining an optimal time step. We note that a similar analysis was done for Trotterization in Ref. [80]
@@ VI Conclusions
In this paper, we have proposed a method to classically optimize BW circuits to efficiently simulate the dynamics of quantum many-body systems, building on the works of Refs. [49, 50, 51, 52]. Like Refs. [51, 52], we make use of MPO approximations of the time evolution operator to variationally update unitary gates in the BW circuits, allowing the method to efficiently scale with system size. The previous works chose particular parameterisations of the two-qubit gates to enforce unitarity, which were optimised using gradient descent approaches. However, our method makes use of the polar decomposition to find efficient and optimal unitary gates [46]. We successively apply our method to multiple Hamiltonians with neighbouring three-body interactions (whereas previous works were limited to just two-body interactions), demonstrating that the dynamics of many-body Hamiltonians can be efficiently implemented using a BW structure with a shallow depth. This is significant because BW circuits can easily be implemented on current digital quantum computers [46].
@@ Acknowledgements.
L.C. was supported by an EPSRC Doctoral prize from the University of Nottingham. A.M. acknowledges the support of the DAAD-WISE scholarship from the Deutscher Akademischer Austausch Dienst, and the KVPY fellowship from the Government of India. At the University of Toronto, A.M. is supported by the Lachlan Gilchrist Fellowship. F.P. acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy–EXC–2111–390814868, the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 851161), as well as the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. A.S. acknowledges support from a research fellowship from the The Royal Commission for the Exhibition of 1851. We acknowledge access to the University of Nottingham Augusta HPC service.
@@
The variational optimisation method proposed in Sec. III worked well for CI in the symmetry protected phase ($g<0$) and had little dependence on initial state. Conversely, the approach can struggle for $g\gtrsim 0$, often getting stuck in local minimum. A reliable strategy to combat this is to anneal the parameter $g$ over the range $[-1,1]$, using the result from a previous choice of $g$ as an initial seed. In practice, we start from a value of $g=-1.0$ and optimise a random initial guess using our variational method. We then iteratively reduce the parameter $g\leftarrow g-0.05$ and optimise using the previous result as the initial guess. This is repeated until the target value of $g$ is obtained. The results of this approach are shown in Fig. 13 for $g=0.25$ and circuit depth $M=3$. The top panel shows the error density and the bottom panel shows the error in the conservation of energy. The results closely match that of Fig. 7.
Figure 13: Error scaling in the trivial phase of the CI model. (Top) The error density, $\delta(\Delta t)$, and (bottom) the error in the conservation of energy, $\mathcal{C}(U_{\rm BW}(\Delta t),\hat{H}_{\rm CI})$ for the CI model with $g=0.25$ and circuit depth $M=3$. All quantities are defined as in Fig. 7.
@@ Appendix B Implementing brickwall and Trotterized circuits on quantum computers
$R_{zz}(\Delta t)=e^{-i\Delta t\hat{Z}\otimes\hat{Z}}$ (27)
can be implemented with a single-qubit rotation gate $R_{z}(\Delta t)=e^{-i\Delta t\hat{Z}}$ placed on the target qubit in-between two identical CNOTs [60]. Notice that this gate is used in the Trotterization for both the CI model and the NNNI model.
Figure 14: Decompositions of unitary gates. (a) A decomposition for any two-qubit unitary gate in terms of single-qubit unitaries and three CNOT gates. (b) The two-qubit rotation gate $R_{zz}(\Delta t)=e^{-i\Delta t\hat{Z}\otimes\hat{Z}}$ can be implemented using only two CNOT gates, and a single-qubit rotation $R_{z}(\Delta t)=e^{-i\Delta t\hat{Z}}$. (c) The three-qubit gate $R_{zxz}(\Delta t)=e^{-i\Delta t\hat{Z}\otimes\hat{X}\otimes\hat{Z}}$ can be decomposed as four CNOT gates, an $R_{z}(\Delta t)$ gate and two Hadamard gates, denoted by $H$. (d) The PXP gate, $R_{pxp}(\Delta t)=e^{-i\Delta\hat{P}\otimes\hat{X}\otimes\hat{P}}$, can be decomposed into eight CNOT gates, six $R_{z}(\pm\Delta t/4)$ rotation gates, two Hadamard gates and four Pauli-X gates, denoted by $X$. (e) A CNOT gate which acts on two next-nearest-neighbouring qubits can be decomposed into 4 CNOTs which act only on neighbouring qubits.
@@
$R_{zxz}(\Delta t)=e^{-i\Delta t\hat{Z}\otimes\hat{X}\otimes\hat{Z}},$ (28)
@@
with $R_{zzz}(\Delta t)=e^{-i\Delta t\hat{Z}\otimes\hat{Z}\otimes\hat{Z}}$. The $R_{zzz}(\Delta t)$ gate can then be decomposed into four CNOTs and an $R_{z}(\Delta t)$ gate [60].
@@
$R_{pxp}(\Delta t)=e^{-i\Delta t\hat{P}\otimes\hat{X}\otimes\hat{P}}\\ =(\hat{X}\otimes H\otimes\hat{X})\,R_{qzq}(\Delta t)\,(\hat{X}\otimes H\otimes\hat{X}),$ (30)
where $\hat{Q}=\hat{\mathds{1}}-\hat{P}$ and $R_{qzq}(\Delta t)=e^{-i\Delta t\hat{Q}\otimes\hat{Z}\otimes\hat{Q}}$. We can then write

### v2

@@ I Introduction
In this paper, we propose a tensor network (TN) approach to classically optimise brickwall (BW) unitary circuits. Our method approximates the time propagator for Hamiltonian dynamics at small times as a matrix product operator (MPO) [53, 54, 55]. This MPO is then used along with the polar decomposition [56, 46] to variationally optimise a unitary BW circuit on a classical computer by maximising their “overlap”. The advantage of using an MPO to approximate the time propagator is that it can reach a desired accuracy with resources which scale only linearly in system size [57, 58, 59]. The classically optimised circuit can thus be used as part of a larger calculation on a digital quantum computer to reach a total time larger than what is accessible on a classical computer.
Our paper is structured as follows. In Sec. II, we give a brief introduction to simulating quantum dynamics with Trotterization. The optimisation procedure is then explained in Sec. III. We demonstrate our approach in Sec. IV for various three-body Hamiltonians. Using numerous metrics, we compare our results to Trotterization to demonstrate the practical advantage of classically optimised BW circuits. We show that the optimised circuits outperform Trotterization in two ways: (i) they approximate the true time evolution operator with less error than Trotterization, and (ii) they can be implemented on digital quantum computers with less two-body entangling layers, thus reducing the error from decoherence. We also benchmark convergence for our method. In Sec. V, we explore the trade-off between optimisation errors and decoherence errors from the device. We use these considerations to explain how to estimate the optimal time step which minimises the compounded error. We give our conclusions in Sec. VI.
@@ III Classical optimisation of unitary circuits
In this section, we explain how one can optimise a BW unitary circuit with two-site gates and depth $M$ to approximate $U(\Delta t)$ (for small $\Delta t$) in a way that is scalable with system size $N$. They are arranged compactly such that each layer has a unitary gate acting on each lattice site (except for lattice sites at the edges), see Figs. 2(a, b) for visualisations of circuit depths $M=3,4$. Note that, while there is no guarantee that a BW circuit with a fixed depth $M$ can well approximate $U(\Delta t)$, the operator entanglement of $U(\Delta t)$ for any local Hamiltonian is bounded by Lieb-Robinson bounds [62] and thus it is reasonable to expect that for sufficiently small times, $U(\Delta t)$ can be well-approximated by a BW circuit with a small depth [63]. Additionally, unlike the Trotter circuits, there is no direct way to determine an optimal BW circuit for general $\hat{H}$. However, they have the distinct advantage that they can be easily implemented on current quantum computers with a number of layers of two-body entangling gates that is constant in $M$, see App. B.
@@ III.1 MPO approximation of $U(\Delta t)$
The application of each gate will give a new MPO with a bond dimension $\chi_{\rm exact}=r\chi$, where $r>1$ is some integer which depends on the circuit. Thus, the bond dimension of the resulting MPO will grow exponentially in the number of applications of the Trotterized circuit. To prevent this, at each time step, we approximate the MPO with another MPO with a smaller bond dimension $\chi^{\prime}$ such that $\chi\leq\chi^{\prime}\leq\chi_{\rm exact}$. In practice, this is achieved by compressing the MPO using SVDs after each application of unitary gate [66]. We allow the bond dimension to grow dynamically, with its value determined through the SVD. We use an SVD truncation error of $\epsilon=10^{-16}$ to ensure that the truncation error is minimal  [^2] See Ref. [66] for a definition of the truncation error.. Notice that this set-up allows for each individual bond dimension between lattice sites to take its own value. Assuming a maximal bond dimension $\chi$, the cost of constructing $U(\Delta t)$ as an MPO is $\mathcal{O}(\Delta t\Delta\tau^{-1}\chi^{3}d^{8})$.
@@ III.4 Contracting the environment
As is often done in TN methods, it is convenient and efficient to recycle partially contracted networks [66]. For this method, it will be useful to save each of the blocks $L_{j}$ and $R_{j}$ to memory for two reasons. The first is the blocks $L_{k-1}$ and $R_{k+1}$ will be the same for each of the $M$ gates which act on site $k$. They can be calculated and then recycled to update each of the $M$ gates. The second reason is that the blocks $L_{j}$ and $R_{j}$ can be used iteratively to calculate the blocks at all sites $j$. That is, $L_{j}$ can be calculated using $L_{j-1}$ and $R_{j}$ can be calculated using $R_{j+1}$. This procedure is illustrated in Fig. 6 for a circuit depth $M=3$, although the exact calculations will differ for other circuit depths. The left column shows how to construct the blocks for even $j$, and the right column for odd $j$. The block $L_{j-1}$ is contracted with all the gates which act on both sites $j-1$ and $j$, along with the adjoint of the MPO tensor at site $j$ to form the block $L_{j}$. Similarly, the block $R_{j+1}$ is contracted with the gates which act on both sites $j$ and $j+1$, along with the adjoint of the MPO tensor at site $j$ to give $R_{j}$. The cost of expanding such blocks is at most $\mathcal{O}\left(\chi^{2}d^{2+2\lceil M/2\rceil}+\chi Md^{4+2\lceil M/2\rceil}\right)$, and storing to memory is at most $\mathcal{O}\left(\chi d^{2\lceil M/2\rceil}\right)$ per block. See Table. 1 for a complete description of all the computational costs in our method.
Estimating $\bm{U(\Delta t)}$ as an MPO
$\mathcal{O}(N\Delta t\,\Delta\tau^{-1}\chi^{3}d^{8})$
Updating the environment
$\mathcal{O}(\chi^{2}d^{2+2\lceil M/2\rceil})+\mathcal{O}(\chi Md^{4+2\lceil M/2\rceil})$
Updating the gates (per gate)
$\mathcal{O}(\chi^{2}d^{2+2\lceil M/2\rceil})+\mathcal{O}(\chi Md^{4+2\lceil M/2\rceil})$
Cost per sweep
    $\mathcal{O}(N\chi^{2}d^{2+2\lceil M/2\rceil})+\mathcal{O}(N\chi Md^{4+2\lceil M/2\rceil})$
Table 1: The computational complexity for each of the steps in the optimisation method. $N$ is the number of lattice sites; $d$ is the local dimension of the system; $\Delta t$ is the time step of unitary $U(\Delta T)=e^{-i\Delta t\hat{H}}$; $\Delta\tau$ is the time step in Trotterization; $\chi$ is the bond dimension of the MPO $U(\Delta t)$; $M$ is the depth of the BW circuit.
@@ IV.1.3 The next-nearest-neighbour Ising model
Unlike the two previous models, we find that the optimisation for the NNNI model is highly susceptible to becoming stuck in local minima. Because of this, optimising the circuits to achieve a scaling that beats $\mathcal{O}(\Delta t)$ is extremely difficult, and only happens on rare occasions for some choice of initial guess. This effect becomes more apparent for larger system sizes. As such, this model serves as an interesting example where the solution is known explicitly, but the variational optimisation method is unable to reliably obtain a good result. Instead, we propose an initial guess which uses the structure of the circuit in Fig. 9(a), but only for the SWAP gates, see Fig. 9(b). All other gates are set to identity. With this choice of initial guess, we find the error metrics converge to those of the second order Trotterization solution, see Figs. 9(c, d) for the error density, and Figs. 9(e, f) for the conservation of energy. These results could motivate the future investigation of strategies to systemically suggest more optimal choices of initial guesses. Notice that in this instance, we observe little improvement when compared to the results of Trotterization.
@@ IV.2 Number of CNOT layers
Table 2: We compare the circuits from Trotterization to ones from our optimisation method. The maximal number of CNOT layers required to implement a single time step of the Trotterized circuits and BW circuits for any system size is shown in the second column. The final column shows the error density $\delta(\Delta t)$ (rounded to three significant figures), for time step $\Delta t=0.1$ and system size $N=8$. Results are shown for Cluster Ising (CI) model, the PXP model and a next-nearest-neighbour Ising (NNNI) model. The order of the approximation, $k$, is given for each circuit. Note that the error density scales as $\mathcal{O}(\Delta t^{k+1})$ for small $\Delta t$. The CI model is for $g=-0.75$ and the NNNI model is for $g_{x}=g_{zz}=g_{z1z}=1$.
@@ V Optimal time step
We have explained how to optimise a BW circuit to efficiently simulate quantum dynamics for some time $\Delta t$, which can then be used repeatedly to achieve some target time $t$ that is unobtainable on a classical computer. However, it is not immediately clear what time step $\Delta t$ should be chosen to simulate the dynamics. On one hand, we have shown that choosing a smaller time step can reduce the algorithmic error for evolving to time $t$. Indeed, if a BW circuit of depth $M$ gives a $k$th order approximation (e.g. $M=3$ for the CI model has order $k=2$), then the cost of evolving to time $t$ is $\mathcal{O}(t\Delta t^{k})$. This error is illustrated by the dashed red line in Fig. 11. This entices us to pick a small time step, resulting in a deeper quantum circuit. On the other hand, quantum computers are never completely isolated from their environment. This introduces errors which are native to the device. We assume that this error is exponential in the circuit depth, as shown by the blue dashed line in Fig. 11. This indicates that our simulations would benefit from a shallower quantum circuit. The total simulation error is a combination of the two, shown by the black curve, and the optimal time step is the one which minimises this error. In what follows, we explain how to estimate each of these errors to guide us in determining an optimal time step. We note that a similar analysis was done for Trotterization in Ref. [82]
@@ VI Conclusions
In this paper, we have proposed a method to classically optimize BW circuits to efficiently simulate the dynamics of quantum many-body systems, building on the works of Refs. [49, 50, 51, 52, 86]. Like Refs. [51, 52], we make use of MPO approximations of the time evolution operator to variationally update unitary gates in the BW circuits, allowing the method to efficiently scale with system size. The previous works chose particular parameterisations of the two-qubit gates to enforce unitarity, which were optimised using gradient descent approaches. However, our method makes use of the polar decomposition to find efficient and optimal unitary gates [46]. We successively apply our method to multiple Hamiltonians with neighbouring three-body interactions (whereas previous works were limited to just two-body interactions), demonstrating that the dynamics of many-body Hamiltonians can be efficiently implemented using a BW structure with a shallow depth. This is significant because BW circuits can easily be implemented on current digital quantum computers [46].
@@ Acknowledgements.
L.C. was supported by an EPSRC Doctoral prize from the University of Nottingham. A.M. acknowledges the support of the DAAD-WISE scholarship from the Deutscher Akademischer Austausch Dienst, and the KVPY fellowship from the Government of India. At the University of Toronto, A.M. is supported by the Lachlan Gilchrist Fellowship. F.P. acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy–EXC–2111–390814868, the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 851161), as well as the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. A.G.-S. acknowledges support from a research fellowship from the The Royal Commission for the Exhibition of 1851. We acknowledge access to the University of Nottingham Augusta HPC service.
@@
The variational optimisation method proposed in Sec. III worked well for CI in the symmetry protected phase ($g<0$) and had little dependence on initial state. Conversely, the approach can struggle for $g\gtrsim 0$, often getting stuck in local minimum. A reliable strategy to combat this is to anneal the parameter $g$ over the range $[-1,1]$, using the result from a previous choice of $g$ as an initial seed. In practice, we start from a value of $g=-1.0$ and optimise a random initial guess using our variational method. We then iteratively reduce the parameter $g\leftarrow g-0.05$ and optimise using the previous result as the initial guess. This is repeated until the target value of $g$ is obtained. The results of this approach are shown in Fig. 13 for $g=0$ and $g=0.25$ with circuit depth $M=3$. The top panel shows the error density and the bottom panel shows the error in the conservation of energy. The results closely match that of Fig. 7.
Figure 13: Error scaling in the CI model. (a, b) The error density, $\delta(\Delta t)$, and (c, d) the error in the conservation of energy, $\mathcal{C}(U_{\rm BW}(\Delta t),\hat{H}_{\rm CI})$ for the CI model with circuit depth $M=3$. The left column is at the tricritical point $g=0$, and the right column is in the trivial phase $g=0.25$. All quantities are defined as in Fig. 7.
@@ Appendix B Implementing brickwall and Trotterized circuits on quantum computers
$R_{zz}(\Delta t)=e^{-i\frac{\Delta t}{2}\hat{Z}\otimes\hat{Z}}$ (27)
can be implemented with a single-qubit rotation gate $R_{z}(\Delta t)=e^{-i\frac{\Delta t}{2}\hat{Z}}$ placed on the target qubit in-between two identical CNOTs [60]. Notice that this gate is used in the Trotterization for both the CI model and the NNNI model.
Figure 14: Decompositions of unitary gates. (a) A decomposition for any two-qubit unitary gate in terms of single-qubit unitaries and three CNOT gates. (b) The two-qubit rotation gate $R_{zz}(\Delta t)=e^{-i\frac{\Delta t}{2}\hat{Z}\otimes\hat{Z}}$ can be implemented using only two CNOT gates, and a single-qubit rotation $R_{z}(\Delta t)=e^{-i\frac{\Delta t}{2}\hat{Z}}$. (c) The three-qubit gate $R_{zxz}(\Delta t)=e^{-i\frac{\Delta t}{2}\hat{Z}\otimes\hat{X}\otimes\hat{Z}}$ can be decomposed as four CNOT gates, an $R_{z}(\Delta t)$ gate and two Hadamard gates, denoted by $H$. (d) The PXP gate, $R_{pxp}(\Delta t)=e^{-i\frac{\Delta t}{2}\hat{P}\otimes\hat{X}\otimes\hat{P}}$, can be decomposed into eight CNOT gates, six $R_{z}(\pm\Delta t/4)$ rotation gates, two Hadamard gates and four Pauli-X gates, denoted by $X$. (e) A CNOT gate which acts on two next-nearest-neighbouring qubits can be decomposed into 4 CNOTs which act only on neighbouring qubits.
@@
$R_{zxz}(\Delta t)=e^{-i\frac{\Delta t}{2}\hat{Z}\otimes\hat{X}\otimes\hat{Z}},$ (28)
@@
with $R_{zzz}(\Delta t)=e^{-i\frac{\Delta t}{2}\hat{Z}\otimes\hat{Z}\otimes\hat{Z}}$. The $R_{zzz}(\Delta t)$ gate can then be decomposed into four CNOTs and an $R_{z}(\Delta t)$ gate [60].
@@
$R_{pxp}(\Delta t)=e^{-i\frac{\Delta t}{2}\hat{P}\otimes\hat{X}\otimes\hat{P}}\\ =(\hat{X}\otimes H\otimes\hat{X})\,R_{qzq}(\Delta t)\,(\hat{X}\otimes H\otimes\hat{X}),$ (30)
where $\hat{Q}=\hat{\mathds{1}}-\hat{P}$ and $R_{qzq}(\Delta t)=e^{-i\frac{\Delta t}{2}\hat{Q}\otimes\hat{Z}\otimes\hat{Q}}$. We can then write